OpenStudy (anonymous):

Consider the equation 5x + 2y = 30 please explain step by step how to graph in slope-intercept form

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

would I take the values from the standard form and plug them in?

OpenStudy (skullpatrol):

The first thing you need to know is the "slope-intercept" form: $$\Huge y=mx+b$$

OpenStudy (skullpatrol):

Can you manipulate the standard form into this form?

OpenStudy (anonymous):

isolate the 2y first

OpenStudy (anonymous):

so subtract the 5x to the other side

OpenStudy (anonymous):

which would be...?

OpenStudy (anonymous):

2y=-5x+30

OpenStudy (anonymous):

Substitute both x and y with 0 in turn, then solve the two equations you come up with. 5(0) + 2y = 30 and 5x + 2(0) = 30

OpenStudy (anonymous):

afterwards divide everything by 2

OpenStudy (anonymous):

to isolate the y

OpenStudy (anonymous):

which would make it y=-5/2x+15

OpenStudy (anonymous):

oh, thank you @chicharito4781 that helped so much

OpenStudy (anonymous):

no problem yo.